# Regularized Estimation of Loading Matrix in Factor Models for High-Dimensional Time Series

### v2

@@ 5.2 Study on $m$
Table 4: Mean and standard deviation (in parentheses) of $|m-\widehat{m}|$ for the example in Section 5.2.
$\delta$ $m$ $p$ $m/3$ varimax1 varimax2 SO sparse
0 $3p^{1/2}$ 50 7 29.1(10.4) 0.9(1.2) 87.9(27.1) 3.3(13.1)
$3p^{2/3}$ 50 14 7.3(6.1) 3.0(1.7) 68.4(23.5) 1.5(7.5)
$3p^{3/4}$ 50 19 16.6(9.5) 5.5(2.5) 72.3(17.3) 0.8(4.6)
$1.8p$ 50 30 18.6(8.4) 14.6(3.5) 50.3(7.7) 2.2(3.4)
$3p^{1/2}$ 100 10 36.8(16.8) 1.6(1.6) 176.7(56.8) 2.9(18.7)
$3p^{2/3}$ 100 22 7.2(8.0) 7.6(2.5) 119.9(51.0) 0.9(8.4)
$3p^{3/4}$ 100 32 0.6(1.6) 15.3(3.4) 133.7(50.4) 0.8(6.2)
$1.8p$ 100 60 15.3(10.1) 47.1(4.9) 99.6(17.0) 5.0(4.6)
$3p^{1/2}$ 200 14 46.5(25.6) 3.0(2.4) 330.1(112.1) 5.0(34.7)
$3p^{2/3}$ 200 34 5.9(7.4) 17.1(3.4) 186.2(104.4) 0.2(0.6)
$3p^{3/4}$ 200 53 0.5(2.3) 36.5(4.9) 222.8(93.1) 0.4(0.9)
$1.8p$ 200 120 6.1(5.4) 139.8(7.7) 199.0(37.9) 8.3(7.2)
$3p^{1/2}$ 500 22 54.7(26.2) 6.7(3.3) 566.0(92.1) 11.3(85.4)
$3p^{2/3}$ 500 63 3.7(7.2) 48.9(5.7) 520.0(297.6) 0.5(0.9)
$3p^{3/4}$ 500 106 3.1(2.3) 113.0(6.8) 438.3(209.4) 0.7(1.4)
$1.8p$ 500 300 60.3(9.0) 542.3(10.5) 507.7(91.2) 13.6(18.0)
0.25 $3p^{1/2}$ 50 7 38.5(11.8) 1.0(1.3) 96.8(24.0) 5.8(16.7)
$3p^{2/3}$ 50 14 15.4(8.9) 3.0(1.6) 81.2(17.9) 3.6(11.5)
$3p^{3/4}$ 50 19 21.0(9.5) 5.4(2.4) 72.4(13.6) 1.8(6.1)
$1.8p$ 50 30 20.2(8.1) 14.6(3.5) 46.4(8.1) 2.6(4.1)
$3p^{1/2}$ 100 10 56.8(22.3) 1.8(2.2) 171.5(50.8) 10.1(34.1)
$3p^{2/3}$ 100 22 19.1(14.2) 7.6(2.5) 180.4(37.7) 2.1(11.1)
$3p^{3/4}$ 100 32 6.4(7.4) 15.3(3.4) 157.1(32.0) 3.9(15.4)
$1.8p$ 100 60 17.3(10.3) 47.2(5.0) 90.2(17.3) 4.2(5.1)
$3p^{1/2}$ 200 14 87.0(45.2) 3.1(2.2) 254.5(67.9) 25.2(82.4)
$3p^{2/3}$ 200 34 17.8(16.2) 17.2(3.4) 375.8(90.4) 1.7(15.3)
$3p^{3/4}$ 200 53 4.8(10.0) 36.4(4.8) 354.6(57.2) 3.4(19.0)
$1.8p$ 200 120 6.5(5.8) 140.1(7.6) 172.6(41.4) 6.7(8.5)
$3p^{1/2}$ 500 22 128.0(89.1) 6.6(3.7) 544.3(33.8) 94.5(245.2)
$3p^{2/3}$ 500 63 12.2(17.1) 49.0(5.7) 631.0(115.5) 11.5(76.2)
$3p^{3/4}$ 500 106 4.6(4.8) 113.2(6.9) 854.2(179.7) 6.9(41.4)
$1.8p$ 500 300 60.2(9.3) 542.1(10.3) 431.0(89.4) 12.9(22.7)
5.3 Orthogonal loadings
We consider a case where ${\mathbf{A}}^{s}$ is a block matrix, which is orthogonal. We can write ${\mathbf{A}}^{s}=\text{{\rm diag}}({\mathbf{a}}_{1},{\mathbf{a}}_{2},{\mathbf{a}}_{3})$, where ${\mathbf{a}}_{1}$ is a $0.4p\times 1$ vector, ${\mathbf{a}}_{2}$ is a $0.3p\times 1$ vector and ${\mathbf{a}}_{3}$ is a $0.3p\times 1$ vector. We report the estimation results in Table 5 and Table 6. It shows that our proposed algorithm still works well when the loading matrix is orthogonal.
Table 5: Mean and standard deviation (in parentheses) of $|m-\widehat{m}|$ for the example in Section 5.3 when the loading matrix is orthogonal.
$\delta=0$ $\delta=0.25$
$n$ $p$ varimax1 varimax2 SO sparse varimax1 varimax2 SO sparse
100 20 27.9(4.2) 3.8(4.0) 26.2(8.3) 3.2(4.0) 30.2(3.6) 6.3(5.0) 27.3(6.3) 4.2(4.5)
100 50 51.6(12.4) 4.3(2.6) 61.0(23.5) 5.1(6.5) 64.3(10.6) 5.4(6.1) 67.0(16.7) 7.8(10.4)
100 100 74.0(23.8) 16.1(3.5) 113.4(46.0) 8.4(13.7) 110.1(21.6) 14.2(6.8) 135.5(29.4) 14.1(22.0)
100 200 87.0(46.9) 54.2(5.8) 222.7(95.9) 13.3(23.3) 178.5(49.6) 52.4(9.1) 268.5(53.0) 30.6(45.8)
100 500 60.0(71.4) 230.0(9.1) 523.6(236.4) 21.6(32.4) 286.4(135.9) 233.1(9.1) 670.8(106.6) 91.6(132.1)
200 20 21.5(5.1) 1.2(1.7) 27.3(8.1) 2.1(3.3) 24.8(4.6) 1.9(2.5) 27.7(7.3) 3.0(4.4)
200 50 32.4(11.8) 4.3(2.1) 64.7(22.9) 3.1(6.4) 46.9(11.3) 4.1(2.3) 68.9(18.1) 5.0(8.7)
200 100 31.5(21.6) 16.9(3.5) 118.3(49.3) 3.4(5.7) 66.9(25.1) 16.7(3.6) 131.3(37.8) 6.0(10.6)
200 200 20.5(24.4) 54.6(5.8) 226.5(106.2) 5.4(10.7) 83.6(45.0) 54.6(5.8) 270.2(79.0) 11.8(25.7)
200 500 15.3(5.8) 229.1(8.4) 543.0(256.0) 8.7(10.8) 64.6(65.1) 229.9(8.4) 709.9(113.9) 26.6(46.4)
500 20 12.2(5.3) 0.5(0.7) 28.7(7.6) 1.6(3.6) 16.0(5.5) 0.5(0.8) 29.6(6.2) 2.4(4.7)
500 50 10.2(7.4) 4.5(2.0) 67.0(25.4) 1.5(5.2) 23.2(10.5) 4.5(1.9) 73.0(16.7) 3.0(8.3)
500 100 4.3(7.4) 16.7(3.7) 129.8(50.1) 1.3(4.6) 22.0(16.8) 16.7(3.7) 142.1(37.6) 3.8(11.7)
500 200 0.9(1.6) 53.8(5.6) 249.1(104.4) 2.0(6.4) 14.7(17.9) 53.9(5.7) 295.5(70.2) 5.3(18.3)
500 500 15.5(3.9) 228.9(8.0) 585.4(276.3) 5.2(13.9) 15.8(7.2) 229.2(8.0) 784.0(115.3) 12.2(41.6)
Table 6: Mean and standard deviation (in parentheses) of the distance between the estimated loading space and the true loading space for the example in Section 5.3 when the loading matrix is orthogonal.
$\delta$ $n$ $p$ eigen varimax1 varimax2 SO sparse
0 100 20 0.129(0.059) 0.128(0.059) 0.108(0.064) 0.227(0.150) 0.098(0.060)
100 50 0.117(0.036) 0.115(0.036) 0.091(0.030) 0.196(0.132) 0.086(0.031)
100 100 0.118(0.032) 0.113(0.034) 0.109(0.019) 0.177(0.122) 0.087(0.029)
100 200 0.119(0.036) 0.107(0.040) 0.157(0.017) 0.172(0.122) 0.088(0.031)
100 500 0.120(0.034) 0.095(0.037) 0.278(0.011) 0.164(0.106) 0.089(0.027)
200 20 0.074(0.029) 0.073(0.029) 0.054(0.030) 0.179(0.140) 0.053(0.028)
200 50 0.071(0.020) 0.067(0.021) 0.062(0.013) 0.127(0.105) 0.049(0.017)
200 100 0.068(0.019) 0.058(0.022) 0.092(0.012) 0.094(0.059) 0.047(0.015)
200 200 0.068(0.018) 0.050(0.019) 0.144(0.010) 0.086(0.044) 0.047(0.014)
200 500 0.067(0.015) 0.045(0.010) 0.270(0.009) 0.081(0.030) 0.046(0.011)
500 20 0.040(0.013) 0.037(0.013) 0.028(0.011) 0.135(0.112) 0.027(0.012)
500 50 0.040(0.009) 0.031(0.010) 0.052(0.011) 0.082(0.062) 0.025(0.008)
500 100 0.038(0.009) 0.025(0.009) 0.084(0.010) 0.060(0.029) 0.024(0.007)
500 200 0.039(0.008) 0.023(0.006) 0.140(0.009) 0.051(0.019) 0.024(0.006)
500 500 0.039(0.008) 0.030(0.004) 0.267(0.008) 0.046(0.014) 0.024(0.005)
0.25 100 20 0.168(0.080) 0.168(0.080) 0.150(0.086) 0.348(0.155) 0.136(0.085)
100 50 0.171(0.061) 0.170(0.061) 0.141(0.064) 0.427(0.152) 0.136(0.062)
100 100 0.187(0.061) 0.184(0.062) 0.155(0.059) 0.496(0.118) 0.150(0.063)
100 200 0.206(0.071) 0.201(0.074) 0.198(0.059) 0.532(0.096) 0.170(0.074)
100 500 0.230(0.071) 0.218(0.077) 0.309(0.031) 0.565(0.055) 0.196(0.072)
200 20 0.096(0.041) 0.095(0.041) 0.073(0.044) 0.297(0.167) 0.073(0.042)
200 50 0.102(0.029) 0.099(0.030) 0.079(0.023) 0.343(0.172) 0.075(0.027)
200 100 0.106(0.029) 0.100(0.031) 0.105(0.016) 0.392(0.174) 0.079(0.025)
200 200 0.115(0.029) 0.102(0.033) 0.155(0.013) 0.474(0.147) 0.088(0.025)
200 500 0.125(0.029) 0.100(0.032) 0.278(0.010) 0.552(0.062) 0.098(0.024)
500 20 0.051(0.017) 0.049(0.018) 0.035(0.015) 0.239(0.149) 0.036(0.017)
500 50 0.056(0.013) 0.051(0.015) 0.057(0.011) 0.263(0.158) 0.038(0.012)
500 100 0.060(0.015) 0.048(0.017) 0.089(0.011) 0.313(0.173) 0.041(0.012)
500 200 0.065(0.014) 0.046(0.015) 0.143(0.010) 0.404(0.164) 0.045(0.012)
500 500 0.073(0.015) 0.050(0.011) 0.270(0.009) 0.536(0.082) 0.053(0.012)
@@ 5.4 Study on sparsity
In this section, we use an example to illustrate the performance of our method on sparsity estimation of the loading matrix. In ${\mathbf{A}}^{s}$, the first $0.4p$ elements in the first column, the middle $0.5p$ elements in the second column, and the last $0.6p$ elements in the third column are nonzero, while all other elements are zero. Note that the second column of ${\mathbf{A}}^{s}$ contains nonzero elements that overlap with those in both the first and third columns. We consider $\delta=0$, $n=200,500$ and $p=20,100$ and $500$. Let $m_{i}$ represent the true number of nonzero elements in ${\bf q}_{i}$ and $\widehat{m}_{i}$ represent the estimated number of nonzero elements in $\widehat{{\mathbf{q}}}_{i}$. In order to evaluate the accuracy of sparity estimation, we also report false negative value (FN, the number of elements falsely identified as zero), false positive (FP, the number of elements falsely identify as nonzero) and F1 score, a number between 0 and 1, which measures the classification accuracy (nonzero or zero). The higher the value of F1 score is, the better the identification is. Table 7 shows the summary results of different measures about identifying non/zero elements for each loading vector. Overall, the proposed approach can identify non/zero elements for each loading vector well. In particular, as sample size increases, the accuracy gets better.
@@ 6 Real data analysis
We apply the proposed algorithm to the Hawaii tourism data and present the results to demonstrate how our proposed method enhances the model interpretability.Hawaii is one of the most popular tourist destinations, attracting over 10 million tourists in 2019. Tourism is the largest single source of Hawaii’s GDP, generating \$17.75 billion in visitor spending. Therefore, a deep understanding of its dynamics is crucial for the local industry and economy (Liu et al.,, 2015).
We downloaded the data from the official website of Hawaii’s government (https://dbedt.hawaii.gov/visitor/). The dataset contains the number of domestic visitors to Hawaii from Washington, D.C., and all 49 U.S. states with $p=50$. The data are recorded monthly from January 2009 to December 2019 with $n=132$. To stabilize the variance and remove the increasing trend, we first took the logarithm transformation and then took a difference to pre-process the original data.
We set $h_{0}=1$ and $r=2$ for demonstration purposes.
Figures 1-3 present the sign of estimated loadings by the method in Lam et al., (2011), by the method in Uematsu and Yamagata, 2022a , and by our method, respectively, where positive loadings are shown in blue, negative loadings in red, and zero loadings in white. The number of zero loadings in the estimate by Lam et al., (2011) is 0, by Uematsu and Yamagata, 2022a is 18, while the number of zero loadings obtained by our method is 41 which is much more sparse. Note that Lam et al., (2011) and Uematsu and Yamagata, 2022a pick the factor with the strongest signal as factor 1, while our approach selects the factor with most sparse loadings as factor 1.
From the results obtained by our method in Figure 3, we can tell that the states with positive loadings on Factor 1 are all located in high latitudes. For Factor 2, most of the states with positive loadings are either inland or in relatively low latitudes, while those with negative loadings have coastlines along the ocean or the Great Lakes. This suggests that Hawaii tourism may be driven by two distinct groups of visitors: those seeking to escape the cold (Factor 1) and those drawn to the beach and water activities (Factor 2).
@@
Figure 1: Estimated loadings of two factors for Hawaii tourism data based on eigen decomposition. Left panel: Estimated loadings for Factor 1; Right panel: Estimated loadings for Factor 2.
Figure 2: Estimated loadings of two factors for Hawaii tourism data based on SO. Left panel: Estimated loadings for Factor 1; Right panel: Estimated loadings for Factor 2.
Figure 3: Estimated loadings of two factors for Hawaii tourism data. Left panel: Estimated loadings for Factor 1; Right panel: Estimated loadings for Factor 2.

### v3

@@ 5.2 Study on $m$
Table 4: Mean and standard deviation (in parentheses) of |mm^||m-$\widehat${m}| for the example in Section 5.2.
$m$ $p$ $m/3$ $\delta=0$ $\delta=0.25$
$3p^{1/2}$ 50 7 3.3(13.1) 5.8(16.7)
$3p^{2/3}$ 50 14 1.5(7.5) 3.6(11.5)
$3p^{3/4}$ 50 19 0.8(4.6) 1.8(6.1)
$1.8p$ 50 30 2.2(3.4) 2.6(4.1)
$3p^{1/2}$ 100 10 2.9(18.7) 10.1(34.1)
$3p^{2/3}$ 100 22 0.9(8.4) 2.1(11.1)
$3p^{3/4}$ 100 32 0.8(6.2) 3.9(15.4)
$1.8p$ 100 60 5.0(4.6) 4.2(5.1)
$3p^{1/2}$ 200 14 5.0(34.7) 25.2(82.4)
$3p^{2/3}$ 200 34 0.2(0.6) 1.7(15.3)
$3p^{3/4}$ 200 53 0.4(0.9) 3.4(19.0)
$1.8p$ 200 120 8.3(7.2) 6.7(8.5)
$3p^{1/2}$ 500 22 11.3(85.4) 94.5(245.2)
$3p^{2/3}$ 500 63 0.5(0.9) 11.5(76.2)
$3p^{3/4}$ 500 106 0.7(1.4) 6.9(41.4)
$1.8p$ 500 300 13.6(18.0) 12.9(22.7)
@@ 5.3 Study on sparsity
In this section, we use an example to illustrate the performance of our method on sparsity estimation of the loading matrix. In ${\mathbf{A}}^{s}$, the first $0.4p$ elements in the first column, the middle $0.5p$ elements in the second column, and the last $0.6p$ elements in the third column are nonzero, while all other elements are zero. Note that the second column of ${\mathbf{A}}^{s}$ contains nonzero elements that overlap with those in both the first and third columns. We consider $\delta=0$, $n=200,500$ and $p=20,100$ and $500$. Let $m_{i}$ represent the true number of nonzero elements in ${\bf q}_{i}$ and $\widehat{m}_{i}$ represent the estimated number of nonzero elements in $\widehat{{\mathbf{q}}}_{i}$. In order to evaluate the accuracy of sparsity estimation, we also report false negative value (FN, the number of elements falsely identified as zero), false positive (FP, the number of elements falsely identify as nonzero) and F1 score, a number between 0 and 1, which measures the classification accuracy (nonzero or zero). The higher the value of F1 score is, the better the identification is. Table 5 shows the summary results of different measures about identifying non/zero elements for each loading vector. Overall, the proposed approach can identify non/zero elements for each loading vector well. In particular, as sample size increases, the accuracy gets better.
@@ 6 Real data analysis
We apply the proposed algorithm to the Hawaii tourism data and present the results to demonstrate how our proposed method enhances the model interpretability. Tourism is the largest single source of Hawaii’s GDP. Therefore, a deep understanding of its dynamics is crucial for the local industry and economy (Liu et al., 2015).
We downloaded the data from the official website of Hawaii’s government (https://dbedt.hawaii.gov/visitor/). The dataset contains the number of domestic visitors to Hawaii from Washington, D.C., and all 49 U.S. states with $p=50$. The data are recorded monthly from January 2009 to December 2019 with $n=132$. To stabilize the variance and remove the increasing trend, we first took the logarithm transformation and then took a difference to pre-process the original data. Set $h_{0}=1$. We use the sequential test proposed in Trapani (2018) to estimate the number of factors and it yields $\widehat{r}=1$. To better compare our methods and others, we present the results with two factors and $\widehat{r}=2$.
First, we use one-step-ahead rolling forecasting to compare different models, including the standard method (Lam et al., 2011), the method in Wu et al. (2026) with the tuning parameter selected based on cross-validation (“SFM (CV)”) and based on BIC (“SFM (BIC)”), and our proposed method. In particular, we use VAR(1) to predict the factor processes. Table 6 reports the mean squared error (MSE) of one-step-ahead rolling forecasting of Hawaii data based on the predictions of the last 30 observations and the number of zeros in the estimated loading matrix based on the full data set. It can be seen that our proposed method produces the most sparse loading matrices and the lowest prediction errors.
Table 6: Hawaii tourism: Comparison of different methods
MSE #of 0
Standard method 1.020 0
SFM (CV) 1.109 0
SFM (BIC) 1.135 16
Our method 1.016 41
Figure 2 plots the estimated loadings for two factors obtained from our method, with loadings for factor 1 in the left panels and loadings for factor 2 in the right panels. In the plot, positive loadings are in blue, negative loadings in red, and zero loadings in white. The top panels show the estimated loadings by a standard method with no sparsity assumption in Lam et al. (2011). None of loadings is zero, and factors are linear combination of all series, making the factors difficult to interpret. The bottom panels show the estimated loadings of our method with 41 zero loadings. We can clearly see the patterns of the factors from the sparse loading matrix. The states that load positively on Factor 1 are all in high latitude. For Factor 2, most of the states that load positively locate inland or are in relatively low latitude, while the states that load negatively have ocean or Great Lakes coastlines. It implies that the tourism of Hawaii may be driven by two groups of visitors — people who would like to escape the cold and who would like to enjoy the beach and water activities.
@@
Figure 2: Estimated loadings for Hawaii tourism data.
The results show that our proposed method captures the dynamics of Hawaii tourism data and improves the prediction accuracy. Moreover, our method yields the most sparse estimate and thus enhances the model interpretability.